# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 5. What to include:
Figure 14: Prompts for the best and worst ad among the initial 64 ads.
Prompt for generating the T-BoN $(T=10)$ ad for GreenBite (plant-based burger patty) "**Key Message:**
*Core Components:**
**Stylistic Qualities:**
Figure 15: Prompts for T-BoN ($T=10$)
**Core Components:**
Figure 16: Prompts for GEPA ($T=10$)
Appendix B Scenarios
Below are ten ad campaign scenarios with creative briefs, outlining the strategic and creative direction for each experimental scenario. The selected campaigns cover a diverse range of products across distinct categories. Each campaign addresses a unique consumer need—whether related to sustainability, luxury, technology, or lifestyle—ensuring comprehensive coverage of various market segments. From environmentally-conscious millennials to affluent, experience-driven travelers, the approach accounts for the full spectrum of consumer interests. By focusing on different consumer needs, the structure enables a holistic marketing strategy that captures both niche and broad demands, providing a well-rounded framework for the study.
B.1 Scenario 1: GreenBite Plant-Based Burger
  • Product: "GreenBite," a new plant-based burger patty.
  • Brand Personality & Tone: Positive, confident, and mouth-watering.
B.2 Scenario 2: AuraSonics X1 Earbuds
  • Brand Personality & Tone: Minimalist, intelligent, and calm.
B.3 Scenario 3: Odyssey E-SUV
  • Product: The "Odyssey E-SUV," a new all-electric family SUV.
B.4 Scenario 4: Oasis Eco-Lodge
  • Brand Personality & Tone: Elegant, peaceful, and understated.
B.5 Scenario 5: Momentum Digital Banking
  • Brand Personality & Tone: Modern, empowering, and simple.
B.6 Scenario 6: MindGarden Meditation App
  • Brand Personality & Tone: Gentle, approachable, and encouraging.
B.7 Scenario 7: Aeterno Watch
  • Brand Personality & Tone: Elegant, timeless, and confident.
B.8 Scenario 8: SyncFlow B2B Software
  • Brand Personality & Tone: Professional, efficient, and innovative.
Appendix C Extended experiment results
T-BoN BO
@@ Appendix C Extended experiment results
Table 2: Progress of T-BoN BO across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline. Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set. The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
@@
Table 3: Progress of GEPA across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline. Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set. The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
@@
T-BoN BO (T=10) Refer to caption Score: 3.107 Refer to caption Score: 3.324 Refer to caption Score: 3.024 Refer to caption Score: 3.257 Refer to caption Score: 3.127

### v3

@@ 5. What to include:
Figure A2: Prompts for the best and worst ad among the initial 64 ads generated using the meta-prompt in Figure A1.
C.3 Prompts used for Critic LLM Operators
comparison_\_prompt =
RESPOND WITH ONLY THE NUMBER 1 OR 2. NOTHING ELSE.
1 = first image is better
2 = second image is better ”
Figure A3: $M_{\text{critic}}$’s pairwise tournament prompt for digital marketing example.
VISUAL EXAMPLES – BEST VS WORST PERFORMING:
RANK 1/10 WORST PERFORMING (Score: 2.14/5.0):
Prompt excerpt: <prompt>…
[Image]
RANK 2/10 LOWER HALF (Score: 2.52/5.0):
Prompt excerpt: <prompt>…
[Image]
$\ldots$ RANK 9/10 UPPER HALF (Score: 3.81/5.0):
Prompt excerpt: <prompt>…
[Image]
RANK 10/10 BEST PERFORMING (Score: 4.15/5.0):
Prompt excerpt: <prompt>…
[Image]
Figure A4: Meta reflection prompt for digital marketing example.
Textual gradient generating meta-prompt for ads (at $t>0$) You are an expert at optimizing image generation prompts for advertising effectiveness. CURRENT PROMPT TO IMPROVE:
{current_prompt}
PERFORMANCE ANALYSIS FROM PREVIOUS ITERATIONS:
{performance_analysis}
TASK: Based on the performance analysis above, generate specific, actionable improvements to make the current prompt more effective. Focus on implementing the successful visual patterns identified in the analysis while avoiding the ineffective elements. Provide 3–5 specific, implementable suggestions for improvement. (It can be addition of a new prompt part, deletion of existing prompt part, or rewriting a prompt part). Each suggestion should reference insights from the performance analysis. RESPONSE FORMAT: 1. [Specific improvement suggestion based on performance analysis] 2. [Specific improvement suggestion based on performance analysis] 3. $\ldots$ Be concrete and actionable. Focus on changes that will meaningfully improve ad effectiveness based on the analysis.
Textual gradient generating meta-prompt for ads (at $t=0$) You are an expert at optimizing image generation prompts for advertising effectiveness. CURRENT PROMPT TO IMPROVE:
{current_prompt}
TASK: Generate specific, actionable improvements to make this ad prompt more effective. Focus on elements that will increase engagement and appeal to the target audience. Consider improvements in: 1. Visual composition and framing 2. Emotional appeal and messaging 3. Color palette and lighting 4. Subject positioning and action 5. Brand integration and logo placement 6. Text overlay space and readability Provide 3–5 specific, implementable suggestions for improvement. Each suggestion should be concrete and actionable. RESPONSE FORMAT: 1. [Specific improvement suggestion based on performance analysis] 2. [Specific improvement suggestion based on performance analysis] 3. $\ldots$ Be concise but specific. Focus on changes that will meaningfully improve ad effectiveness.
Figure A5: Meta prompts for improving ad image-generation prompts (with and without performance analysis).
Meta-prompt for applying textual gradients You are an expert at revising image generation prompts based on improvement suggestions. ORIGINAL PROMPT:
{current_prompt}
IMPROVEMENT SUGGESTIONS:
{gradient}
TASK: Rewrite the prompt incorporating the improvement suggestions while maintaining the core message and structure. REQUIREMENTS: Keep the prompt structure and format similar to the original Integrate the improvement suggestions naturally Maintain coherence and readability Ensure the prompt is optimized for image generation Keep the prompt length reasonable (not too long) OUTPUT CONSTRAINT:
Return ONLY the revised prompt, no explanations or additional text.
Figure A6: Meta-prompt for applying improvement suggestions to revise an image-generation ad prompt.
C.4 Prompt for Evaluation Based on Persona
Persona prompt for simulating ad effectiveness. SYSTEM: $\{$
You are an AI assistant. Your task is to answer the TASK as if you are the individual described in the ‘Persona Profile’ (which contains their past survey responses). Remain consistent with the persona’s past survey responses and stated characteristics. Carefully follow any instructions provided for the new question, including formatting requirements.
PERSONA DATA: $\{$
Which part of the United States do you currently live in?
Question Type: Single Choice
Options:
1 - Northeast (PA, NY, NJ, RI, CT, MA, VT, NH, ME)
2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
3 - South (TX, OK, AR, LA, KY, TN, MS, AL, WV, DC, MD, DE, VA, NC, SC, GA, FL)
4 - West (WA, OR, ID, MT, WY, CA, NV, UT, CO, AZ, NM)
5 - Pacific (HI, AK)
Answer: 2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
What is the highest level of schooling or degree that you have completed?
Question Type: Single Choice
Options:
1 - Less than high school
2 - High school graduate
3 - Some college, no degree
4 - Associate’s degree
5 - College graduate/some postgrad
6 - Postgraduate
Answer: 3 - Some college, no degree
$\ldots$ (Many other survey questions and answers) $\ldots$
Suppose you were given \$5 and had to offer to another (anonymous) person a way to split the money. The other person can either accept or reject your offer. If the other person accepts your offer, you would each receive the amount you proposed. If the other person rejects your offer, you would both receive \$0. How much would you offer to the other person?
Question Type: Single Choice
Options:
1 - \$0
2 - \$1
3 - \$2
4 - \$3
5 - \$4
6 - \$5
Answer: 3 - \$2
$\ldots$ (Many other survey questions and answers) $\ldots$
Effective Score Scale Definition:
No explanation. Just the score.
Figure A7: Prompt for simulating the effectiveness of a given ad-persona combination.
Appendix D Appendix for Section 8
D.1 Implementation Details of GEPA for digital ad optimization
Implementation overview
GEPA [agrawal2025gepa] is an iterative, prompt-optimization-based, self-improving AI method. The key idea of GEPA is to “combine reflection with evolution”. Here, an evolution is the same concept as taking a textual gradient in textgrad, and reflection is the same reflection TextBO (\$6) uses. GEPA iterates over three steps in each optimization iteration: Step 1) choosing a prompt and evaluating it; Step 2) updating the meta-reflection; and Step 3) improving the prompt.
GEPA optimizes the prompt for a set of problem instances. In standard agentic benchmarks considered in agrawal2025gepa, a problem instance is a single task in a benchmark (together with whatever information is needed to evaluate it) that can be stored and revisited. In our digital ad setting, a problem instance corresponds to an individual ad viewer (a persona): each evaluation asks how effective a given ad is for a specific ad viewer (i.e., an ad–ad viewer combination), and the overall score for an ad is computed by averaging effectiveness ratings over a random sample of ad viewers (personas).
Unlike the agentic benchmarks, in ad optimization, ad platforms do not have the ability to store and retrieve a particular ad viewer at will. Instead, ad platforms (or managers) test an ad to a population of ad viewers and receive an aggregated evaluation signal (e.g., an average effectiveness rating across a sampled set of personas), which prevents us from maintaining instance-level records, such as which prompts win on which ad viewer (in our case, a persona). Therefore, to apply GEPA for digital ad optimization, we need to adapt it.
For Step 1), in agentic AI experiments considered in §9 and agrawal2025gepa), where we can store, retrieve, or evaluate individual problem instances), it uses instance-level information to (i) maintain a Pareto archive of which prompts in the candidate prompt set currently win on which problem instances, (ii) stochastically sample a non-dominated candidate prompt (favoring candidates that win on more instances), and (iii) evaluate that sampled candidate by rolling it out on a minibatch of problem instances. On the other hand, in the digital ad optimization experiment (§7), where algorithms have access only to aggregated evaluation results, this step’s procedure simplifies to that of TextBO: evaluate the prompt across the entire test persona set.
In Step 2), we employ the same meta-reflection procedure as used for TextBO.
For Step 3), we take one step of evolution (or equivalently, one step of textual gradient) of the prompt we evaluated in Step 1) and accept it only if it has a better evaluation than the original prompt. For agentic AI experiments considered in $\S$9 and agrawal2025gepa) where we have access to individual problem instances, when the offspring prompt is accepted, we append it to the candidate prompt set and don’t throw away the original prompt from the candidate set. On the other hand, in digital ad optimization in $\S$7, where we cannot maintain the instance-level Pareto archive and thus cannot construct the frequency-weighted, non-dominated sampling distribution, we discard the original prompt from the candidate set. This is because, if we were to keep appending accepted prompts, the candidate prompt set would grow without bound, and the evaluation budget per prompt would be uniformly diluted. Therefore, when the offspring prompt is accepted, we discard the original prompt, keeping the candidate prompt set size constant.
Notably, after adapting GEPA to the digital ad optimization problem setup, where we cannot store and retrieve personas freely, it effectively reduces to TextBO with $G=1$ and $N=1$; we describe the pseudocode of the adapted version of GEPA in Algorithm 3.
Shared components with TextBO.
To ensure a fair comparison, we keep the non-algorithmic components identical to those used for TextBO (§ 7.2): (i) the same scenario-specific initial pool of 64 prompts generated from the creative brief, (ii) the same ad-generation module $\Phi$ (Imagen 4), (iii) the same critic model $M_{\text{critic}}$ (Gemini 2.5 Flash) and the same meta-prompts for reflection and prompt rewriting, and (iv) the same persona-based evaluation operator $\mathrm{Eval}(\cdot)$ induced by Twin-2k-500 personas and LLM-scored effectiveness ratings.
Candidate prompt pool (“population”).
In the ad setting, each prompt $\pi$ produces a single ad image $\Phi(\pi)$ and a single scalar score $s=\mathrm{Eval}(\Phi(\pi),\mathcal{D}_{\mathcal{X}})$. As we described earlier, we implement GEPA with a fixed-size candidate pool (population) of size $J=5$. We initialize this pool using the same Worst5-of-64 prompts described in $\S$7.3. For each candidate $j\in\{1,\dots,J\}$ we maintain a triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ where $s_{t}^{j}$ is the empirical evaluation score.
Evaluation
Each time we evaluate an ad image $\Phi(\pi)$ during training, we follow the same protocol as in § 7.2: we randomly sample 200 personas from the training persona set and compute the mean effectiveness score across those personas (using the same 1–5 rubric and the same logprob-to-expectation conversion). This yields the scalar feedback used by GEPA for selection. For reporting final performance (plots/tables), we evaluate selected ads on the held-out test personas.
Meta-reflection update.
After collecting evaluations, GEPA constructs a global meta-reflection $R_{t}$ using the same MetaReflect operator as TextBO (§ 6 and § 7.2). Concretely, the critic is prompted to compare a subset of the best- and worst-performing creatives observed so far (subject to context-length limits) and to distill recurring patterns into a small set of actionable rules (e.g., “prefer warm lighting,” “show social context,” “avoid clutter”). This reflection is then used to guide subsequent prompt evolution steps.
Evolution (mutation) step and acceptance rule.
Given a current candidate prompt $\pi_{t}^{j}$ and reflection $R_{t}$, we generate a single offspring prompt by (i) sampling a targeted textual improvement using the critic (a TextGrad-style local edit [textgrad]) and (ii) rewriting the prompt accordingly:
$\delta_{t}^{j}\sim\nabla_{\text{text}}(\pi_{t}^{j};R_{t},M_{\text{critic}}),\qquad\tilde{\pi}_{t}^{j}=\operatorname{Apply}(\pi_{t}^{j},\delta_{t}^{j}).$
We then generate the corresponding offspring creative $\Phi(\tilde{\pi}_{t}^{j})$ and evaluate it to obtain $\tilde{s}_{t}^{j}=\mathrm{Eval}(\Phi(\tilde{\pi}_{t}^{j}),\mathcal{D}_{\mathcal{X}})$. We use an elitist acceptance rule: if $\tilde{s}_{t}^{j}>s_{t}^{j}$, the offspring replaces the parent in the pool; otherwise the parent is kept. This keeps the pool size constant across iterations (preventing unbounded growth of candidates under a fixed evaluation budget) while ensuring that each candidate lineage is non-decreasing in its accepted scores.
Full procedure and hyperparameters.
As in TextBO, we run GEPA for $T=10$ iterations with pool size $J=5$ per scenario. At each iteration, we evolve each of the $J$ candidates once (one offspring per candidate), evaluate the offspring using 200 sampled training personas, apply the acceptance rule above, and then update the shared meta-reflection. For comparison plots and tables, we track the best-performing candidate in the pool at each iteration, $j_{t}^{\star}:=\arg\max_{j}s_{t}^{j}$, and report the corresponding ad $\Phi(\pi_{t}^{j_{t}^{\star}})$.
Input: System $\Phi$; initial prompts $\{\pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; iterations $T$; pool size $J$
Output: Best ads $\{\Phi(\pi_{t}^{j_{t}^{\star}})\}_{t=0}^{T}$, where $j_{t}^{\star}=\arg\max_{j}s_{t}^{j}$
1 for $j=1$ to $J$ do
2 $c_{0}^{j}\leftarrow\Phi(\pi_{0}^{j})$
3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}})$
[…]
5
6for $t=0$ to T1T-1 do
7 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}})$
8 for $j=1$ to $J$ do
9    $\delta_{t}^{j}\leftarrow\nabla_{\text{text}}(\pi_{t}^{j};R_{t},M_{\text{critic}})$
10    $\tilde{\pi}_{t+1}^{j}\leftarrow\operatorname{Apply}(\pi_{t}^{j},\delta_{t}^{j})$
11    $\tilde{c}_{t+1}^{j}\leftarrow\Phi(\tilde{\pi}_{t+1}^{j})$
12    $\tilde{s}_{t+1}^{j}\leftarrow\text{Eval}(\tilde{c}_{t+1}^{j},\mathcal{D}_{\mathcal{X}})$
13    if $\tilde{s}_{t+1}^{j}>s_{t}^{j}$ then
14       $(\pi_{t+1}^{j},c_{t+1}^{j},s_{t+1}^{j})\leftarrow(\tilde{\pi}_{t+1}^{j},\tilde{c}_{t+1}^{j},\tilde{s}_{t+1}^{j})$
15
16    else
17       $(\pi_{t+1}^{j},c_{t+1}^{j},s_{t+1}^{j})\leftarrow(\pi_{t}^{j},c_{t}^{j},s_{t}^{j})$
18
19 $H_{t+1}\leftarrow H_{t}\cup\{(\pi_{t+1}^{j},c_{t+1}^{j},s_{t+1}^{j})\}_{j=1}^{J}$
20
return $\{\Phi(\pi_{t}^{j_{t}^{\star}})\}_{t=0}^{T}$
Algorithm 3 GEPA (adapted for ad optimization)
D.2 Detailed Experiment Results
TextBO
@@ D.2 Detailed Experiment Results
Table A1: Progress of TextBO across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline. Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set. The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
@@
Table A2: Progress of GEPA across 10 optimization steps from its starting point (Worst5-of-N) and comparison with the Best-of-N baseline. Each data point, in the mean (standard error) format, indicates the mean and standard error across 411 personas in the test persona set. The last column, labeled “Average”, reports the averages of the means across the eight scenarios.
@@
TextBO (T=10) Refer to caption Score: 3.107 Refer to caption Score: 3.324 Refer to caption Score: 3.024 Refer to caption Score: 3.257 Refer to caption Score: 3.127